OpenStudy (anonymous):

If the measure of DEF is 76, what is the measure of DE?

OpenStudy (anonymous):

@jim_thompson5910 @Mertsj @Jonask

jimthompson5910 (jim_thompson5910):

is this the wrong picture?

OpenStudy (anonymous):

no :)

jimthompson5910 (jim_thompson5910):

you mention DEF, but points D, E, or F are not in the picture

OpenStudy (anonymous):

OH omygosh Iam sooo sorry xD I didn't mean to put that lol

jimthompson5910 (jim_thompson5910):

that's fine, so the actual problem is just what we see in the pic?

OpenStudy (anonymous):

Yeah!! lol @jim_thompson5910 that's the correct one

OpenStudy (anonymous):

I subtracted then divided

jimthompson5910 (jim_thompson5910):

we're using the same formula from a few problems ago (m greater arc XZ - m lesser arc XZ)/2 = angle XYZ (230 - q)/2 = 50 230 - q = 50*2 230 - q = 100 q = ???

OpenStudy (anonymous):

I got 130?

jimthompson5910 (jim_thompson5910):

same here

OpenStudy (sirm3d):

another way to look at the problem is that 1 complete circle measures 360 degrees. the major arc is 230 degrees, so the minor arc is what's left from 360 degrees.

OpenStudy (anonymous):

Okay thank you so much!! I have another simular question to this coming up :)